A 30 kg mass is resting on a rough, horizontal surface. A horizontal force of 100 N is required just to start it moving. What is the approximate value for the coefficient of static friction?
1
Expert's answer
2020-12-14T07:30:56-0500

"F_s=\mu_s N"


where "F_s" =friction force

"N" = normal force

"\mu_s" = coefficient of static friction


which can be written as;


"\frac{F}{mg}=\mu"


where F= force

m= mass

g= gravity

"\mu=" coefficient of friction


substituting the given values in the equation


"\frac{100}{30\times10}=\frac{1}{3}"


Coefficient of static friction = 1/3